I have to rewrite the middle term as the sum of two terms and then factor by grouping.
x^2 + 16x + 60

Answer by mszlmb(115)

x^2 + 16x + 60

what two numbers add up to 16 and multiply to 60?
60=30*2=15*4=5*12
60=3*20=6*10 aha!  difference of 16

so x2+6x+10x+60  factor like terms
x(x+6)+10(x+6)  put like terms together
(x+10)(x+6)
;)